# Towards Adversarial Robustness via Debiased High-Confidence Logit Alignment

### v1

@@ Abstract
Despite the significant advances that deep neural networks (DNNs) have achieved in various visual tasks, they still exhibit vulnerability to adversarial examples, leading to serious security concerns. Recent adversarial training techniques have utilized inverse adversarial attacks to generate high-confidence examples, aiming to align the distributions of adversarial examples with the high-confidence regions of their corresponding classes. However, in this paper, our investigation reveals that high-confidence outputs under inverse adversarial attacks are correlated with biased feature activation. Specifically, training with inverse adversarial examples causes the model’s attention to shift towards background features, introducing a spurious correlation bias. To address this bias, we propose Debiased High-Confidence Adversarial Training (DHAT), a novel approach that not only aligns the logits of adversarial examples with debiased high-confidence logits obtained from inverse adversarial examples, but also restores the model’s attention to its normal state by enhancing foreground logit orthogonality. Extensive experiments demonstrate that DHAT achieves state-of-the-art performance and exhibits robust generalization capabilities across various vision datasets. Additionally, DHAT can seamlessly integrate with existing advanced adversarial training techniques for improving the performance.
Introduction
@@
Deep neural networks (DNNs) have demonstrated exceptional performance across a range of visual tasks (Zeng et al. 2024; Li et al. 2024a; Gao et al. 2024). However, these networks remain susceptible to adversarial attacks, which present significant security challenges (Szegedy et al. 2014; Fawzi, Fawzi, and Fawzi 2018). Adversarial attacks introduce subtle perturbations that are imperceptible to humans but can significantly disrupt the inference process of DNNs, leading to incorrect predictions. Adversarial training (AT) is widely recognized as one of the most effective methods for defending against these attacks by incorporating adversarial examples into the training process (Goodfellow, Shlens, and Szegedy 2014a; Shafahi et al. 2019).
@@
Recently, numerous advanced AT-based techniques have been proposed to address the vulnerabilities of DNNs (Li and Liu 2023; Subramanian et al. 2024; Huang et al. 2023). Prominent methods in this domain focus on aligning the distributions of adversarial examples with regions corresponding to their true classes, thereby enhancing classification accuracy under adversarial attacks. For example, MART (Wang et al. 2020) and TRADES (Zhang et al. 2019) aim to align the logits of adversarial examples with their corresponding natural examples, promoting consistency in predictions. Expanding on this, UIAT (Dong et al. 2023) and ACR (Cho, Han, and Kim 2023) extend this approach by generating inverse adversarial examples (Salman et al. 2021) that exhibit higher confidence compared to their natural counterparts. This strategy helps alleviate the negative impact of misclassification of natural examples in alignment and directs adversarial examples toward regions with correct high-confidence classifications.
Figure 1: Impact of inverse adversarial examples on model attention. (a) Attention maps generated by Grad-CAM (Selvaraju et al. 2017a) under adversarial and inverse adversarial attacks using the traditional AT model. (b) Comparison of IoU(%) between model attention regions and the true foreground and background regions using UIAT and AT.
Upon investigation, we find that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation. By visualizing the feature attention maps of the conventional AT model under both adversarial attack and inverse adversarial attack conditions (see Figure 1(a)), we observe that adversarial attacks disrupt the model’s usual attention, leading to incorrect predictions. In contrast, inverse adversarial attacks amplify the model’s attention on features, and stronger inverse attacks can further improve prediction accuracy (Dong et al. 2023). Nevertheless, this amplified attention often causes a shift from the foreground to background features.
To further investigate the impact of inverse adversarial examples on adversarial training, we compared the Intersection over Union (IoU) between the model’s attention maps and the foreground/background regions for the UIAT model (Dong et al. 2023) (trained with inverse adversarial examples) as well as the conventional AT model. As shown in Figure 1(b), increasing the magnitude $\epsilon$ of inverse adversarial attacks enhances the model’s attention to image features. However, this enhanced attention to features introduces a bias, with the UIAT model showing a remarkable shift towards emphasizing background features compared to the conventional AT model. This suggests a compelling perspective: AT model tends to focus on background features when using inverse adversarial examples for training, resulting in biased feature activation. This shift introduces spurious correlation bias, where the model over-relies on background features but neglects foreground features, leading to degraded generalization (Han et al. 2023; Zhang et al. 2024).
@@
Building upon the above findings, we introduce a method named Debiased High-Confidence Adversarial Training (DHAT). DHAT implements two key techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and recalibrates the biased high-confidence logits derived from inverse adversarial examples. This regularization aligns the logits of adversarial examples with debiased high-confidence logits to improve adversarial robustness and mitigate spurious correlation bias. FLOE further refines this process by reducing the correlation between high-confidence logits and background features in affine space, which helps restore the model’s attention to its normal state. Extensive experiments demonstrate that DHAT outperforms state-of-the-art techniques across various datasets, providing superior adversarial robustness and generalization. Additionally, our DHAT can be seamlessly integrated with existing advanced adversarial training methods.
@@
  • We identify that the adversarial training model shows a biased feature activation under inverse adversarial attacks. Training with inverse adversarial examples causes the model’s attention to shift towards background features, resulting in spurious correlation bias.
@@
  • Extensive experiments validate the superior robustness and generalization of our method against adversarial attacks. Additionally, our approach integrates seamlessly with existing advanced adversarial training strategies.
Related Work
Adversarial Attack
@@
Numerous studies have investigated the vulnerability of DNNs and have proposed various adversarial attack methods. Goodfellow, Shlens, and Szegedy (2014b) introduced the Fast Gradient Sign Method (FGSM), which compute the adversarial perturbation in the direction of the gradient:
@@
where $\eta$ denotes the attack step size. Building on this, Madry et al. (2018) proposed Projected Gradient Descent (PGD), which iteratively updates the gradient of the input data to generate more potent adversarial examples, thereby increasing the success rate of attacks:
@@
where $\Pi_{\mathbb{B}(\epsilon)}$ projects the perturbation onto the $p$-norm ball of radius $\epsilon$. Furthermore, Carlini and Wagner (2017) developed the Carlini-Wagner (C&W) attack, an optimization-based method capable of producing more imperceptible adversarial examples. Croce and Hein (2020b) introduced AutoAttack, an ensemble attack strategy that is user-independent and parameter-free.
Adversarial Defense
Adversarial training (AT) is a widely used strategy for improving model robustness. The core principle of AT is to include adversarial examples in the training process, thereby aligning the model’s outputs towards the true label distribution (Xie et al. 2020; Fowl et al. 2021). Mathematically, conventional AT methods are formulated as a min-max optimization problem:
@@
Recent advancements in AT have introduced several techniques to further improve adversarial robustness. Wu, Xia, and Wang (2020) introduced Adversarial Weight Perturbation (AWP), which enhances robustness by flattening the loss landscape with respect to input variations. Kim et al. (2023) proposed Feature Separation and Recalibration (FSR) to capture useful logits from non-robust features. Wei et al. (2023) developed Class-wise Calibrated Fair Adversarial Training (CFA), which adaptively configures training for each class to achieve category-level robust fairness. Li et al. (2024b) proposed SGLR, which uses self-distillation to refine the soft-label distribution logits, thereby calibrating the adversarial training.
Methods similar to our approach include MART (Wang et al. 2020) and UIAT (Dong et al. 2023). These methods use Kullback–Leibler (KL) divergence to align the logits of adversarial examples with high-confidence logits. MART encourages sharing a similar logit distribution between adversarial examples and their corresponding natural examples. UIAT generates inverse adversarial examples (Salman et al. 2021) that exhibit higher confidence compared to natural examples, thereby enhancing the alignment of adversarial examples towards the high-confidence regions of the decision surface. However, our analysis reveals that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation. Specifically, training with inverse adversarial examples causes model to shift attention from the foreground to background features. Consequently, we focus on mitigating spurious correlation bias in robust model trained with inverse adversarial examples.
Methodology
In this section, we first introduce the basic notation. Then, we describe the Debiased High-Confidence Logit Regularization (DHLR). Next, we present the Foreground Logit Orthogonal Enhancement (FLOE). Finally, we detail the optimization strategy of our proposed method.
Notation
We denote the training dataset as $(X,Y)=\{(x_{i},y_{i})|~{}i\in(1,2,\dots,N)\}$, where $x$ and $y$ denote the natural input examples and their corresponding labels, respectively. $N$ represents the total number of examples in the dataset. Let $\theta$ denote the parameters of a convolutional neural network (CNN) for image classification. The network’s output logits are denoted as $f_{\theta}(\cdot)$, and the final predicted label is given by $\arg\max f_{\theta}(\cdot)$. Inverse adversarial examples $\check{x}$ can be regarded as examples for which the model correctly predicts the label with high-confidence. These examples are generated by applying image perturbations aimed at minimizing the attack loss (Salman et al. 2021):
@@
Debiased High-Confidence Logit Regularization
@@
First, we quantify the model’s biased activation degree towards background features under inverse adversarial attacks. Following the approach proposed by  (Dong et al. 2020), we use Grad-CAM (Selvaraju et al. 2017b) to generate pixel-wise attention maps $M\in[0,1]^{n\times n}$ for the natural example $x$. We separate the background feature $\check{x}_{(S)}$ from the inverse adversarial examples $\check{x}$ based on the $M$:
$[\check{x}_{(S)}]_{(i,j)}=\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)},$ (6)
where $\mathbb{I}_{(M_{i,j}<\omega)}$ is an indicator function and $\omega$ is a predefined threshold. Next, we compute the logits for the background feature $\check{x}_{(S)}$:
$\check{z}_{(S)}=f_{\theta}(\check{x}_{(S)}).$ (7)
@@
$\check{z}^{*}=\check{z}-\check{z}_{(S)},$ (8)
@@
Foreground Logit Orthogonal Enhancement
Post-hoc recalibration of high-confidence logits through DHLR is still insufficient to restore the model’s attention to its normal state when confronted with inverse adversarial examples. Therefore, we further propose a Foreground Logit Orthogonal Enhancement (FLOE) to mitigate this issue.
The primary motivation behind FLOE is to enhance the model’s focus to foreground features under inverse adversarial attacks. FLOE achieves this by minimizing the correlation of the high-confidence logits $\check{z}$ with respect to the background feature logits $\check{z}_{(S)}$, which is implemented by reducing the projection of $\check{z}$ onto $\check{z}_{(S)}$ in the affine space. The optimization objective for FLOE is formulated as:
$\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)})=|\check{z}-\frac{\check{z}\cdot\check{z}_{(S)}}{|\check{z}_{(S)}|^{2}}\cdot\check{z}_{(S)}|_{p},$ (10)
@@
Optimization Strategy
Finally, we integrate the Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE) techniques into a unified training framework Debiased High-Confidence Adversarial Training (DHAT). The overall optimization objective for DHAT is formulated as:
@@
$+\lambda_{2}\cdot\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)}),$
@@
Experiments
Experimental Settings
Datasets and models.
We conduct experiments on the CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and Imagenette (Howard 2019) datasets. The architecture for our experiments include the mainstream models WRN28-10 (Zagoruyko and Komodakis 2016), ResNet-18 (He et al. 2016), ResNet-50 (He et al. 2016), Inception-V3 (Szegedy et al. 2015), and VGG16 (Simonyan and Zisserman 2014). Detailed information about the datasets is provided in the supplementary material.
Training setup.
We employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$. Each model is trained for 100 epochs, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs. For adversarial example generation, we use a fixed 10-iteration attack. The maximum $\ell_{\infty}$-norm of the adversarial perturbation is set to $\epsilon=8/255$, while the inverse adversarial perturbation is set to $\epsilon=4/255$. Both perturbations use a fixed step size of $\alpha=2/255$. Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.0 across all datasets. Additional training details are available in the supplementary material.
Evaluation setup.
For robustness performance evaluation, we employ PGD (Madry et al. 2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) attacks under the $\ell_{\infty}$ norm. AutoAttack comprises APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al. 2020). “Clean” denotes natural examples without adversarial perturbations. Furthermore, we introduce the concept of the robust generalization gap (Yin et al. 2023), referred to as the “Robust Gap”. More details are provided in the supplementary material.

### v2

@@ Abstract
Despite the remarkable progress of deep neural networks (DNNs) in various visual tasks, their vulnerability to adversarial examples raises significant security concerns. Recent adversarial training methods leverage inverse adversarial attacks to generate high-confidence examples, aiming to align adversarial distributions with high-confidence class regions. However, our investigation reveals that under inverse adversarial attacks, high-confidence outputs are influenced by biased feature activations, causing models to rely on background features that lack a causal relationship with the labels. This spurious correlation bias leads to overfitting irrelevant background features during adversarial training, thereby degrading the model’s robust performance and generalization capabilities. To address this issue, we propose Debiased High-Confidence Adversarial Training (DHAT), a novel approach that aligns adversarial logits with debiased high-confidence logits and restores proper attention by enhancing foreground logit orthogonality. Extensive experiments demonstrate that DHAT achieves state-of-the-art robustness on both CIFAR and ImageNet-1K benchmarks, while significantly improving generalization by mitigating the feature bias inherent in inverse adversarial training approaches. Code is available at https://github.com/KejiaZhang-Robust/DHAT.
1 Introduction
@@
Deep neural networks (DNNs) have demonstrated exceptional performance across a range of visual tasks [53, 26, 37, 13]. However, these networks remain susceptible to adversarial attacks, which present significant security challenges [42, 11]. Adversarial attacks introduce subtle perturbations that are imperceptible to humans but can significantly disrupt the inference process of DNNs, leading to incorrect predictions. Adversarial training (AT) is widely recognized as one of the most effective methods for defending against these attacks by incorporating adversarial examples into the training process [14, 36].
Figure 1: Inverse adversarial attack exploits spurious correlations via attention misallocation to non-causal background features. The negative $\epsilon$ denotes inverse adversarial perturbation strength. (a) Grad-CAM [34] visualizations show that adversarial attacks disrupt attention and cause errors, while inverse adversarial attacks amplify activation and improve accuracy but expose a critical flaw: attention shifts from foreground to irrelevant background regions. (b) Quantitative comparison of IoU(%) between attention regions and the true foreground and background regions using UIAT [9] and AT. Increasing inverse attack strength $-\epsilon$ demonstrates UIAT’s disproportionate focus on non-causal background features, exposing reliance on spurious correlations [3, 35, 30].
@@
Recently, numerous advanced AT-based techniques have been proposed to address the vulnerabilities of DNNs [24, 40, 19, 25]. Prominent methods in this domain focus on aligning the distributions of adversarial examples to their true classes, thereby enhancing classification accuracy under adversarial attacks. For example, MART [45] and TRADES [54] aim to align the logits of adversarial examples with their corresponding natural examples, promoting consistency in predictions. Expanding on this, UIAT [9] and ACR [5] extend this approach by generating inverse adversarial examples [32] that exhibit higher confidence compared to their natural counterparts. This strategy helps alleviate the negative impact of misclassification of natural examples in alignment and directs adversarial examples toward regions with correct high-confidence classifications.
Upon investigation, we uncover a remarkable yet concerning phenomenon: inverse adversarial examples produce high-confidence outputs by triggering biased feature activation. To illuminate this effect, we visualize attention maps from a conventional adversarial training (AT) model under both adversarial and inverse adversarial attacks (see Figure 1 (a)). Adversarial attacks disrupt normal attention patterns, causing incorrect predictions. In contrast, inverse adversarial attacks amplify the model’s attention on features, with stronger attacks even improving prediction accuracy [9, 48]. However, this apparent enhancement masks a critical flaw: the model’s attention systematically shifts from discriminative foreground features to irrelevant background regions—akin to identifying a sheep by detecting grass rather than the animal itself. This reliance on spurious correlations undermines the model’s causal reasoning capabilities and compromises its robustness [20, 6, 1].
To delve deeper into this issue, we conduct a statistical analysis using the Intersection over Union (IoU) metric to measure the alignment between attention maps and foreground/background regions on a subset of ImageNet [18]. Leveraging the Class Activation Mapping (CAM)-based technique inspired by [10], we define foreground as regions with high-class activation (i.e., areas critical for predictions) and background as regions with low activation. Comparing the UIAT model [9] (trained with inverse adversarial examples) to a standard AT model reveals a concerning trend. As evidenced in Figure 1 (b), while increasing the magnitude $\epsilon$ of inverse adversarial attacks enhances overall feature attention, the UIAT model fails to improve IoU for foreground features compared to the AT model. Instead, it demonstrates a marked increase in IoU for background features—a clear indicator of attention misallocation.
This leads to a pivotal insight: the inverse adversarial training disproportionately skews attention toward background features, resulting in biased feature activation. This shift engenders a fundamental vulnerability: a spurious correlation bias where the model becomes overly dependent on contextual background cues rather than intrinsic object features. Much like a classifier that associates beaches with birds rather than wings and beaks, this misplaced reliance significantly degrades robustness and generalization performance [15, 55], as conclusively demonstrated in Table 5.
@@
Building upon the above findings, we introduce a method named Debiased High-Confidence Adversarial Training (DHAT). DHAT implements two key techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and debiase the spurious correlations by recalibrating the biased high-confidence logits derived from inverse adversarial examples. This regularization aligns the logits of adversarial examples with debiased high-confidence logits to improve adversarial robustness and mitigate spurious correlation bias. FLOE further refines this process by reducing the correlation between high-confidence logits and background features in affine space, which helps restore the model’s attention to its normal state. Extensive experiments demonstrate that DHAT outperforms state-of-the-art techniques across various datasets, providing superior adversarial robustness and generalization. Additionally, our DHAT can be seamlessly integrated with existing advanced adversarial training methods.
@@
  • We identify that the adversarial training model shows a biased feature activation under inverse adversarial attacks. Training with inverse adversarial examples causes the model’s attention shifting towards background features, thus lead to spurious correlation bias.
@@
  • Extensive experiments validate the superior robustness and generalization of our method against adversarial attacks. Additionally, our approach integrates smoothly with existing advanced adversarial training strategies.
2 Related Work
2.1 Adversarial Attack
@@
Numerous studies have investigated the vulnerability of DNNs and have proposed various adversarial attack methods. Goodfellow et al. [14] introduced the Fast Gradient Sign Method (FGSM), which compute the adversarial perturbation in the direction of the gradient:
@@
where $\eta$ denotes the attack step size. Building on this, Madry et al. [29] proposed Projected Gradient Descent (PGD), which iteratively updates the gradient of the input data to generate more potent adversarial examples, thereby increasing the success rate of attacks:
@@
where $\Pi_{\mathbb{B}(\epsilon)}$ projects the perturbation onto the $p$-norm ball of radius $\epsilon$. Furthermore, Carlini and Wagner [4] developed the Carlini-Wagner (C&W) attack, an optimization-based method capable of producing more imperceptible adversarial examples. Croce and Hein [7] introduced AutoAttack, an ensemble attack strategy that is user-independent and parameter-free.
2.2 Adversarial Defense
Adversarial training (AT) is a widely used strategy for improving model robustness. The core principle of AT is to include adversarial examples in the training process, thereby aligning the model’s outputs towards the true label distribution [50, 12]. Mathematically, conventional AT methods are formulated as a min-max optimization problem:
@@ 2.2 Adversarial Defense
Recent advancements in AT have introduced several techniques to further improve adversarial robustness. Wu et al. [49] introduced Adversarial Weight Perturbation (AWP), which enhances robustness by flattening the loss landscape with respect to input variations. Kim et al. [21] proposed Feature Separation and Recalibration (FSR) to capture useful logits from non-robust features. Wei et al. [47] developed Class-wise Calibrated Fair Adversarial Training (CFA), which adaptively configures training for each class to achieve category-level robust fairness. Li et al. [27] proposed SGLR, which uses self-distillation to refine the soft-label distribution logits, thereby calibrating the adversarial training.
Methods similar to our approach include MART [45] and UIAT [9]. These methods use Kullback–Leibler (KL) divergence to align the logits of adversarial examples with high-confidence logits. MART encourages sharing a similar logit distribution between adversarial examples and their corresponding natural examples. UIAT generates inverse adversarial examples [32] that exhibit higher confidence compared to natural examples, thereby enhancing the alignment of adversarial examples towards the high-confidence regions of the decision surface.
3 Methodology
In this section, we introduce the basic notation and present our approach to mitigate adversarial bias and improve robustness. Our method recalibrates high-confidence logits and reduces reliance on spurious background features. The optimization strategy is outlined, with pseudocode provided in Section 7 of the supplementary material.
3.1 Notation
We denote the training dataset as $(X,Y)=\{(x_{i},y_{i})|~{}i\in(1,2,\dots,N)\}$, where $x$ and $y$ denote the natural input examples and their corresponding labels, respectively. $N$ represents the total number of examples in the dataset. Let $\theta$ denote the parameters of a convolutional neural network (CNN) for image classification. The network’s output logits are denoted as $f_{\theta}(\cdot)$, and the final predicted label is given by $\arg\max f_{\theta}(\cdot)$. Inverse adversarial examples $\check{x}$ can be regarded as examples for which the model correctly predicts the label with high-confidence. These examples are generated by applying image perturbations aimed at minimizing the attack loss [32]:
@@ 3.1 Notation
3.2 Debiased High-Confidence Logit Regularization
@@
First, we quantify the model’s biased activation degree towards background features under inverse adversarial attacks. Following the approach proposed by [10], we use using Grad-CAM [33] or alternative methods such as SAM [22] to generate pixel-wise attention maps $M\in[0,1]^{n\times n}$ for the natural example $x$. We separate the background feature $\check{x}_{(B)}$ from the inverse adversarial examples $\check{x}$ based on the $M$:
$[\check{x}_{(B)}]_{(i,j)}=\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)},$ (6)
where $\mathbb{I}_{(M_{i,j}<\omega)}$ is an indicator function and $\omega$ is a predefined threshold. Next, we compute the logits for the background feature $\check{x}_{(B)}$:
$\check{z}_{(B)}=f_{\theta}(\check{x}_{(B)}).$ (7)
@@
$\check{z}^{*}=\check{z}-\check{z}_{(B)},$ (8)
@@
3.3 Foreground Logit Orthogonal Enhancement
While DHLR effectively recalibrates high-confidence logits to mitigate spurious correlations, it fails to fully restore the model’s focus on foreground features when handling inverse adversarial examples $\check{x}$. This is because DHLR primarily aligns logits with debiased high-confidence targets, without directly addressing the model’s persistent bias toward background features. Under inverse adversarial attacks, the model still exhibits strong activation in irrelevant background regions, undermining its ability to prioritize critical foreground cues.
To address this issue, we introduce Foreground Logit Orthogonal Enhancement (FLOE), a targeted method designed to minimize the correlation between high-confidence logits $\check{z}$ and background feature logits $\check{z}_{(B)}$. The intuition behind FLOE is straightforward: by reducing the projection of $\check{z}$ onto $\check{z}_{(B)}$ in the affine space, we make $\check{z}$ less explainable by $\check{z}_{(B)}$, thereby encouraging the model to rely more on foreground features. This projection-based approach ensures that the model’s predictions are driven by meaningful foreground information rather than spurious background correlations. The optimization objective is formulated as:
$\mathcal{L}_{FLOE}(\check{z},\check{z}_{(B)})=-|\check{z}-\frac{\check{z}\cdot\check{z}_{(B)}}{|\check{z}_{(B)}|^{2}}\cdot\check{z}_{(B)}|_{p},$ (10)
@@ 3.3 Foreground Logit Orthogonal Enhancement
3.4 Optimization Strategy
Finally, we integrate the DHLR and FLOE techniques into a unified training framework Debiased High-Confidence Adversarial Training (DHAT). The overall optimization objective for DHAT is formulated as:
@@ 3.4 Optimization Strategy
$+\lambda_{2}\cdot\mathcal{L}_{FLOE}(\check{z},\check{z}_{(B)}),$
@@
4 Experiments